What is the pH value of each of the following solutions: a) 0.0045 M HCl; b) 0.00683 M NaOH?


1
Expert's answer
2021-01-15T06:26:41-0500

Hcl is strong acid and dissociates completely. So [H+] = 0.0045

Now PH = -log [H+] = -log 0.0045 = 2.35 (ans)

Similarly NaOH is strong base and [OH-] = 0.00683

So POH = -log[OH-] = -log 0.00683 = 2.16

We know that POH + PH = 14

So PH = 14 - POH = 14 - 2.16 = 11.84 (ans)